what is the 288th term of the series a,bb,ccc,dddd.....
Thanks in anticipation .my exam is on 20th Sunday. do ans. soon

Answer by drk(1908)

Sorry I came across this later than you hoped. But just for fun, here is the answer. Notice the pattern 1, 2, 3, 4, . . . We want to sum the integers from 1 to n until we find where 288 lies.
SUM(n=1 to k) n = k(k+1)/2.
k(k+1)/2 = 288
K^2 + K - 576 = 0
Using quadratic, K ~ 23.5.
SO the 288th term should occur in the 24th letter, or X.